OpenStudy (anonymous):

find the slope of the line that contains the points (5,3) and (-2,-2)

OpenStudy (anonymous):

5/7

OpenStudy (anonymous):

slope = difference in y cood / diff in x coods = 3-(-2) / 5-(-2) = 5/7

OpenStudy (anonymous):

\[\frac{-2-3}{-2-5}=\frac{-5}{-7}=\frac{5}{7}\] or \[\frac{3+2}{5+2}=\frac{5}{7}\]